OpenStudy (anonymous):

x²=3x-1 Quadratic Formula can you help me please :)

OpenStudy (solomonzelman):

\[x^2=3x-1\]\[x^2\color{red} {-3x+1}=3x-1\color{red} {-3x+1}\]\[x^2-3x+1=0\]

OpenStudy (solomonzelman):

\[\huge\color{red}{\frac{-b±\sqrt{b^2-4ac}}{2a}}\]

OpenStudy (solomonzelman):

\[\huge\color{red}{1}x^2\huge\color{red}{-3}x+\huge\color{red}{1}=0\]compare to \[\huge\color{red}{a}x^2+\huge\color{red}{b}x+\huge\color{red}{c}=0\]Do you see your a b and c? PLUG THEM IN.

OpenStudy (anonymous):

Thansks :)

OpenStudy (solomonzelman):

\[\huge\color{green}{\frac{-(-3)±\sqrt{(-3)^2-4(1)(1)}}{2(1)}}\]

OpenStudy (isaiah.feynman):

Lol do you have like an advanced version of openstudy how do you do this big, colored text? @SolomonZelman

OpenStudy (anonymous):

*Thanks :)

OpenStudy (solomonzelman):

It should be hard, go to http://openstudy.com/study#/updates/51fbcbade4b0cc46c14a461d I combined some latex, that's all.

OpenStudy (isaiah.feynman):

Thank you.

OpenStudy (solomonzelman):

Anytime!